# When Do LLMs Admit Their Mistakes? Understanding the Role of Model Belief in Retraction

### v2

@@ B.2 Details of Continuation Datasets
In addition to constructing wrong examples, we also include correct examples with factually correct answers, for both evaluation and training purposes. To avoid introducing bias during supervised fine-tuning, we aim to maintain a balanced number of correct and wrong examples. Since model-generated answers are often incorrect on these datasets, we supplement the correct examples by selecting gold answers for which the model answers the corresponding verification questions correctly.
@@
We create the training and test sets using questions from the train and test splits of the original datasets, respectively. However, Qwen2.5-7B and Olmo2-7B know little about the correct answers on Celebrity. As a result, in Table 1, only their test sets for Celebrity are imbalanced, containing 1000 incorrect examples and a smaller number of correct ones. Additionally, these two models do not have a training set for Celebrity, which only impacts the in-distribution supervised fine-tuning experiments discussed in Section 6. Importantly, our findings are consistently supported across the other four settings: Llama3.1-8B + Wikidata, Llama3.1-8B + Celebrity, Qwen2.5-7B + Wikidata, and Olmo2-7B + Wikidata.
B.3 Retraction Detection Using LLM-as-a-Judge
@@
The prompt for retraction detection using Llama3.3-70B-Instruct is shown below. Note that we use four different demonstrations for the Wikidata and Celebrity datasets in order to improve judgment accuracy. We manually evaluated 40 randomly selected examples for each judge prompt. Llama3.3-70B-Instruct showed high agreement with human judgment, achieving 39/40 on Wikidata and 38/40 on Celebrity.
@@
Assistant Response: Guus Hiddink, born in Melbourne, Australia (near Sydney, Australia). Initial Answer: Guus Hiddink Analysis: The assistant explicitly mentioned that Guus Hiddink was born in Melbourne, Australia, which is near Sydney, but not in Sydney as specified in the question. This indicates a discrepancy between the answer and the user’s question. Output: True [Omit the other two examples here…] Now, process the following: User Question: {question} Assistant Response: {response} Initial Answer: {model’s answer}
B.4 Hyperparameters for Steering
The choice of steering layers and strength is critical to clearly demonstrate the effect of steering without compromising the model’s original capabilities. Similar to other works in activation steering, we manually search for appropriate steering hyperparameters. Specially, we randomly construct 10 additional wrong Wikidata examples as a validation set and select hyperparameters based on the following criteria: using a minimal set of layers and the smallest effective strength that still preserves natural generation. Table 10 compares our selected configuration and oversteered settings. Although hyperparameters are chosen using only wrong Wikidata examples for negative belief steering, they generalize well to positive belief steering, positive examples, the Celebrity dataset, and the is-appended setting, demonstrating the generalizability of the belief steering. The final choices are listed in Table 11.
Question Properly Steered Response Oversteered Response
Model Layers Strength $\alpha$
Llama3.1-8B 6-14 1.2
Qwen2.5-7B 10-18 2.5
Olmo2-7B 8-30 1.5
Table 11: Steering hyperparameters.
B.5 Implementation Details for Patching
Patching Attention Weights.
Patching Attention Value Vectors.
B.6 Hyperparameters for Supervised Fine-Tuning
We fine-tune models using LoRA for 2 epochs with a learning rate of $1\text{e}{-4}$ and a batch size of 8, implemented via LLaMA-Factory (Zheng et al., 2024). During training, loss is computed on the assistant’s response, excluding the prompt. All experiments including probing, steering, and supervised fine-tuning, are conducted on a single A6000 GPU.
Appendix C Additional Results
C.1 Probing Plots
Since the retraction recall of Qwen2.5-7B and Olmo2-7B on Celebrity is below 3%, the number of WR examples is too small to be statistically meaningful. Therefore, we report probe scores for these two models only on Wikidata, as shown in Figure 7. Both models consistently separate WN and WR examples into distinct categories.
@@ C.1 Probing Plots
Figure 7: Average probe scores across layers for Qwen2.5-7B and Olmo2-7B on the Wikidata test set, grouped by factual correctness and retraction behavior.
C.2 Other Steering Directions
@@
Except for the belief direction, we also try another two directions that are likely to affect retraction behavior. (1) Wikidata retraction direction: The positive examples are those that the model actually retracts from the Wikidata training set, and negative examples are those that the model does not retract. (2) Wikidata correctness direction: The positive examples contain factually correct answers from the Wikidata training set, and negative examples contain factually incorrect answers. We search for the best hyperparameters as described in Appendix B.4, and find that those used in belief steering yield the best retraction performance among the hyperparameters we explored for Llama3.1-8B. We show the results in Table 12.
@@
C.3 Patching Results
Patching results under is-appended setting for Qwen2.5-7B and Olmo2-7B are shown in Table 16 and  16. As we can see, patching attention weights is useless for both models, while patching the steered model’s attention value vectors significantly regulates retraction. Note that for Olmo2-7B, we increase the original $\alpha$ from $1.5$ to $5.0$ to make belief steering effective under is-appended setting. This implies that, at $\alpha=1.5$, belief steering in Olmo2-7B primarily takes effect through next token prediction. Nevertheless, larger $\alpha$ values still modify the attention value vectors in a manner consistent with our overall conclusions. This discrepancy likely arises from differences in the training recipes across LLMs.
@@ C.3 Patching Results
C.4 Supervised Fine-tuning Results
C.4.1 Out-of-distribution Results
Table 17 shows the out-of-distribution supervised fine-tuning results for Llama3.1-8B. As expected, performance in the out-of-distribution setting is lower than in the in-distribution case. To support broader applicability, developing a diverse, large-scale, retraction-focused supervised fine-tuning dataset holds great promise.
Wikidata Celebrity
Precision Recall Precision Recall
Baseline 0.9012 0.2529 0.7722 0.1477
In-distribution SFT 0.7815 0.8453 0.8988 0.9031
Out-of-distribution SFT 0.5180 0.6705 0.6635 0.3390
UTQA SFT 0.6653 0.5225 0.6244 0.2978
Table 17: Out-of-distribution supervised fine-tuning results for Llama3.1-8B. The first row indicates the evaluation dataset. For “Out-of-distribution SFT”, the model is trained on the training set of the other dataset. “UTQA SFT” denotes fine-tuning on the UTQA training set.
C.4.2 SFT Results for Qwen and Olmo
Building on Llama3.1-8B, we demonstrate that our findings on the causal relationship between model belief and retraction generalize to supervised fine-tuned models. This is further supported by results from Qwen2.5-7B and Olmo2-7B. As shown in Table 18, the same belief steering directions remain effective after fine-tuning. Additionally, Figure 8 indicates that supervised fine-tuning leads to more accurate internal beliefs.
@@ C.4.2 SFT Results for Qwen and Olmo
Precision Recall Precision Recall
@@
(a) Qwen2.5-7B on Wikidata.
(b) Olmo2-7B on Wikidata.
Figure 8: Average probe scores across layers for Qwen2.5-7B, Olmo2-7B (Base), and their fine-tuned variants (SFT).
C.4.3 Practical Application
The continuation setting is a synthetic setup designed to facilitate controlled study. Here, we consider a more realistic scenario: given a question from Wikidata or Celebrity, what does supervised fine-tuning achieve? The results are shown in Table 19. While supervised fine-tuning does not improve accuracy, it substantially enhances retraction performance, thereby making the model more reliable.
Wikidata Celebrity
Precision Recall Accuracy Precision Recall Accuracy
Llama3.1-8B 0.9928 0.2715 0.0841 0.8125 0.0884 0.3163
Llama3.1-8B SFT 0.9481 0.7079 0.0840 0.9774 0.8162 0.2261
Table 19: Supervised fine-tuning results for Llama3.1-8B in a realistic setting. Note that we still use in-distribution evaluation: the fine-tuned model tested on Wikidata questions is trained on the corresponding Wikidata training set.

### v3

@@ A.2 Details of Continuation Datasets
In addition to constructing wrong examples, we also include correct examples with factually correct answers, to evaluate over-retraction and support in-distribution SFT in Section 6. To avoid bias during SFT, we aim to balance the number of correct and wrong examples. Because model-generated answers are often incorrect on these two datasets, we supplement the correct examples by selecting gold answers for which the model answers the corresponding verification questions correctly.
@@
We build the training and test sets using questions from the train and test splits of the original datasets, respectively. However, Qwen2.5-7B and Olmo2-7B know little about the correct answers in Celebrity. Consequently, as shown in Table 1, their Celebrity test sets are imbalanced, containing 1,000 incorrect examples and a smaller number of correct ones. Additionally, these two models lack a Celebrity training set, which only impacts the in-distribution SFT experiments in Section 6. Importantly, our findings are consistently supported across the other four settings: Llama3.1-8B + Wikidata, Llama3.1-8B + Celebrity, Qwen2.5-7B + Wikidata, and Olmo2-7B + Wikidata.
Layers Strength $\alpha$
Llama3.1-8B 6-14 1.2
Qwen2.5-7B 10-18 2.5
Olmo2-7B 8-30 1.5
Table 9: Steering hyperparameters.
A.3 Hyperparameters for Steering
The choice of steering layers and strength is critical to clearly demonstrate the effect of steering without compromising the model’s original capabilities. Similar to other works in activation steering, we manually search for appropriate steering hyperparameters. Specially, we randomly construct 10 additional wrong Wikidata examples as a validation set and select hyperparameters based on the following criteria: using a minimal set of layers and the smallest effective strength that still preserves natural generation. Table 8 compares our selected configuration and oversteered settings. Although hyperparameters are chosen using only wrong Wikidata examples for negative belief steering, they generalize well to positive belief steering, positive examples, the Celebrity dataset, and the is-appended setting, demonstrating the generalizability of the belief steering. While more exhaustive hyperparameter sweeps could potentially identify even better settings, we find that the current choices produce stable and consistent effects across different datasets. The final choices are listed in Table 9.
A.4 Implementation Details for Patching
Patching Attention Weights.
Patching Attention Value Vectors.
A.5 Implementation Details for Supervised Fine-Tuning
For each of our datasets, we synthetically construct an in-distribution supervised fine-tuning training set (e.g., training on Wikidata training set and evaluating on Wikidata test set). Specifically, we append “is the correct answer.” to correct examples that contain factually correct answers in the training dataset, and “is not the correct answer.” to wrong examples.
We fine-tune models using LoRA (Hu et al., 2022) for 2 epochs with a learning rate of $1\text{e}{-4}$ and a batch size of 8, implemented via LLaMA-Factory (Zheng et al., 2024). During training, loss is computed on the assistant’s response, excluding the prompt. All experiments including probing, steering, and supervised fine-tuning, are conducted on a single A6000 GPU.
The results are shown in Table 6 and 20. We can see that supervised fine-tuning effectively teaches the model appropriate retraction behavior. The model learns to distinguish between factually correct and incorrect answers and respond accordingly, i.e., saying “is the correct answer” to correct answers and saying “is not the correct answer” to incorrect ones.
A.6 Retraction Detection Using LLM-as-a-Judge
@@
The prompt for retraction detection using Llama3.3-70B-Instruct [^5] https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct is shown below. Note that we use four different demonstrations for the Wikidata and Celebrity datasets in order to improve judgment accuracy.
To validate judge reliability, we first manually evaluated 40 randomly selected examples for each judge prompt. Llama3.3-70B-Instruct shows high agreement with human judgment, achieving 39/40 on Wikidata and 38/40 on Celebrity.
Additionally, we use GPT-4.1-mini [^6] https://platform.openai.com/docs/models/gpt-4.1-mini to evaluate Llama3.1-8B’s retraction performance under no-steering, positive-steering, and negative-steering settings ($1,202\times 3=3,606$ examples on Wikidata; $826\times 3=2,478$ examples on Celebirty). We report both raw agreement (consistency) and Cohen’s $\kappa$ (Cohen, 1960) between GPT-4.1-mini and Llama3.3-70B-Instruct in Table 10. These results indicate that Llama3.3-70B-Instruct achieves very high agreement with both human annotators and an independent LLM judge.
Consistency Cohen’s $\kappa$
Wikidata 0.9750 0.9445
Celebrity 0.9730 0.9364
Table 10: Judgment consistency between GPT-4.1-mini and Llama3.3-70B-Instruct on retraction evaluation.
@@
Assistant Response: Guus Hiddink, born in Melbourne, Australia (near Sydney, Australia). Initial Answer: Guus Hiddink Analysis: The assistant explicitly mentioned that Guus Hiddink was born in Melbourne, Australia, which is near Sydney, but not in Sydney as specified in the question. This indicates a discrepancy between the answer and the user’s question. Output: True [Omit the other two examples here…] Now, process the following: User Question: {question}
Assistant Response: {response}
Initial Answer: {model’s answer}
@@
Figure 5: Layer-wise AUROC of belief scores for factual correctness and retraction of Qwen2.5-7B and Olmo2-7B on the Wikidata test set.
Figure 6: Layer-wise AUROC of belief scores for factual correctness and retraction of Llama3.1-70B-Instruct on Wikidata.
Appendix B Additional Results
B.1 Probing Plots
Since the retraction recall of Qwen2.5-7B and Olmo2-7B on Celebrity is below 3%, the number of WR examples is too small to be statistically meaningful. Therefore, we report AUROC of belief scores for these two models only on Wikidata, as shown in Figure 5. Both models consistently present high correlation between belief scores and retraction.
Robustness to Model Scale.
To evaluate whetehr our findings generalize to larger LLMs, we additionaly study Llama3.1-70B-Instruct. We construct its Wikidata continuation test set of 4,492 examples with balanced correct and wrong answers, and test its retraction behavior. The model achieves a retraction precision of 0.9126 and recall of 0.5303, outperforming the smaller Llama3.1-8B-Instruct (precision 0.9012, recall 0.2579), yet still leaves substantial room for improvement.
We next examine whether belief continues to explain retraction behavior at this scale through the probing experiment. Given that Llama3.1-70B-Instruct contains 80 layers with 8192-dimensional hidden states (vs. 32 layers and 4096 dimensions in Llama3.1-8B-Instruct), to mitigate overfitting, we expand the UTQA training set from 2,400 to 8,000 examples by additionally incorporating samples from ANLI (Nie et al., 2020), AG News (Zhang et al., 2015), MRPC (Wang et al., 2019), SQuAD (Rajpurkar et al., 2016), OpenBookQA (Mihaylov et al., 2018), Winogrande (Sakaguchi et al., 2020), and Truthful QA (Lin et al., 2022). We train an independent single-linear probe at each layer, and the resulting AUROC scores are shown in Figure 6.
We can see that the belief probes achieve moderately strong performance on predicting retraction behavior in layers 29-35, and consistently outperform factual-correctness prediction, mirroring patterns observed in the 8B variant and in other model families. Their predictive power is somewhat lower than probes on smaller models, possibly because larger models are more expressive and encode multiple entangled features at the coarse layer level. More fine-grained extraction of belief representations, such as at the head level or using sparse autoencoders (Huben et al., 2024; Shu et al., 2025), may reveal more generalizable belief signals.
Overall, these results indicate that the retraction mechanism and its connection to belief remain consistent from smaller to larger LLMs across families.
B.2 Uncertainty vs. Retraction
Figure 7: AUROC of different predictors for retraction in Llama3.1-8B on the Wikidata test set.
It is important to note that “low belief scores” indicate that the model internally regards an answer as incorrect, rather than being uncertain about its correctness. Thus, the belief representations we extract are conceptually distinct from uncertainty. Nonethless, one might expect retraction to occur when a model is uncertain about its answer. In this section, we examine this relationship and report the AUROC of uncertainty scores against retraction labels. All experiments are conducted using LLaMA3.1-8B on the Wikidata dataset.
Token-Level Entropy.
We first examine whether higher uncertainty in an answer, as measured by token-level entropy, is associated with a greater likelihood of retraction. For each answer in the continuation dataset, we compute the average token-level entropy as follows:
$\text{Entropy}=-\frac{1}{T}\sum_{t=1}^{T}\sum_{v\in V}p_{t}(v)\log p_{t}(v),$
where $T$ is the number of tokens in the answer, $V$ is the vocabulary, and $p_{t}(v)$ is the model’s predicted probability of token $v$ at position $t$.
Using token-level entropy to predict retraction yields an AUROC of 0.518, only marginally above random chance ($\text{AUROC}=0.5$). This result indicates no meaningful correlation with retraction.
Consistency-Based Uncertainty.
Next, we assess consistency-based uncertainty (Xiong et al., 2024) by sampling $n=5$ answers per question and defining an answer’s uncertainty as:
$\text{Uncertainty}(a_{i})=1-\frac{|a_{i}|}{n},$
where $|a_{i}|$ is the number of times the same answer appears among the five samples.
Predicting retraction using consistency-based uncertainty yields an AUROC of 0.533, reflecting weak discriminative capacity.
Inter-Answer Entropy.
Additionally, we investigate whether uncertainty can identify questions where the model is more likely to exhibit retraction behavior. We measure question-level uncertainty by computing the entropy over the five generated answers per question, following (Kuhn et al., 2023; Xiong et al., 2024):
$\text{Entropy}(q)=-\sum_{a\in A_{q}}p(a)\log p(a),$
where $A_{q}$ is the set of unique answers generated for question $q$, and $p(a)$ is the relative frequency of answer $a$ among the five samples. To handle semantic equivalence, we extract answers (e.g., person names in the Wikidata dataset) from model responses using Llama-3.3-70B-Instruct, rather than relying on an NLI model as in Kuhn et al. (2023).
Inter-answer entropy achieves an AUROC of 0.505 for predicting retraction, offering little predictive values.
Compared to all the methods discussed above, our belief probe scores show a significantly higher correlation with retraction behavior, as presented in Figure 7. Uncertainty, by contrast, may require more precise definitions and further study to uncover its potential connection to retraction.
B.3 Robustness to Prompting and Decoding Variation
To assess the robustness of belief-retraction dynamics, we evaluate how prompt templates and decoding hyperparameters affect both the model’s baseline retraction behavior and the effectiveness of belief steering. All experiments in this section use Llama3.1-8B-Instruct and Wikidata.
B.3.1 Robustness to Prompt Variation
We test two standard prompt phrasings along with an adversarial variant designed to introduce an incorrect statement before the question to potentially bias the model’s belief:
  • Original Prompt: Name an association football player who was born in Naples, Italy.
  • Prompt 1: Can you name an association football player who was born in Naples, Italy?
  • Prompt 2: Which association football player was born in Naples, Italy? Just name one.
  • Adv Prompt: Barack Obama is a politician who was born in New York City, United States. Name an association football player who was born in Naples, Italy.
Precision Recall
Prompt 1
No Steering 0.8994 0.4609
Belief- Steering 0.5296 0.8935
Belief+ Steering 0.9412 0.0266
Prompt 2
No Steering 0.9211 0.0582
Belief- Steering 0.5115 0.8902
Belief+ Steering 1.0000 0.0017
Adv Prompt
No Steering 0.8138 0.4725
Belief- Steering 0.5280 0.8319
Belief+ Steering 0.9091 0.0166
Table 11: Steering results under different prompt variants.
As shown in Table 11, across all templates, belief- steering greatly leads to more retraction, while belief+ steering suppresses retraction. This demonstrates that belief steering is robust to changes in prompt phrasing.
B.3.2 Robustness to Decoding Variation
We further analyze the effect of decoding hyperparameters. In addtion to greedy decoding, we evaluate temperature sampling ($\text{temperature}=0.7,\text{top-p}=0.95$), repeating each run three times across different seeds.
Precision Recall
GD 0.9012 0.2579
TS 0.8814(0.0105) 0.3128(0.0088)
TS + Belief– 0.5257(0.0013) 0.8980(0.0101)
TS + Belief+ 0.9551(0.0024) 0.0355(0.0019)
Table 12: Steering results under different decoding hyperparameters. Subscripts indicate standard deviation. “GD” denotes Greedy Decoding, and “TS” denotes Temperature Sampling.
Table 12 suggests that the link between belief and retraction holds consistently across decoding hyperparameters.
B.4 Other Steering Directions
@@
Except for the belief direction, we also try another two directions that are likely to affect retraction behavior. (1) Wikidata retraction direction: The positive examples are those that the model actually retracts from the Wikidata training set, and negative examples are those that the model does not retract. (2) Wikidata correctness direction: The positive examples contain factually correct answers from the Wikidata training set, and negative examples contain factually incorrect answers. We search for the best hyperparameters as described in Appendix A.3, and find that those used in belief steering yield the best retraction performance among the hyperparameters we explored for Llama3.1-8B. We show the results in Table 13.
@@
Llama3.1-8B Qwen2.5-7B Olmo2-7B
Wikidata Celebrity Wikidata Celebrity Wikidata Celebrity
[…]
Table 16: Change in attention weights to the answer span.
@@
B.5 Changes in Attention Weights
As shown in Table 16, negative belief steering increases the model’s attention to the entity name when generating the next token, while positive belief steering decreases it. This suggests that belief steering indeed changes attention weights.
B.6 Patching Results
Wikidata Celebrity
Prec. Rec. Prec. Rec.
No Steer 0.9012 0.2579 0.7722 0.1477
belief-
Patch W 0.8325 0.2729 0.7113 0.1671
Patch V 0.5249 0.5441 0.6351 0.3245
Full Steer 0.5157 0.9268 0.4803 0.7676
belief+
Patch W 0.8984 0.1913 0.7333 0.1065
Patch V 0.9700 0.1614 0.6552 0.0920
Full Steer 1.0000 0.0067 0.5217 0.0291
Table 17: Patching results for Llama3.1-8B on continuation test sets.
@@ B.6 Patching Results
Patching results without appending is is shown in Table 17. Patching results under is-appended setting for Qwen2.5-7B and Olmo2-7B are shown in Table 18 and  19. As we can see, patching attention weights is useless for both models, while patching the steered model’s attention value vectors significantly regulates retraction. Note that for Olmo2-7B, we increase the original $\alpha$ from $1.5$ to $5.0$ to make belief steering effective under is-appended setting. This implies that, at $\alpha=1.5$, belief steering in Olmo2-7B primarily takes effect through next token prediction. Nevertheless, larger $\alpha$ values still modify the attention value vectors in a manner consistent with our overall conclusions. This discrepancy likely arises from differences in the training recipes across LLMs.
B.7 Supervised Fine-tuning Results
Building on Llama3.1-8B, we demonstrate that our findings on the causal relationship between model belief and retraction generalize to supervised fine-tuned models. This is further supported by results from Qwen2.5-7B and Olmo2-7B. As shown in Table 20, the same belief steering directions remain effective after fine-tuning. Additionally, Figure 8 and 9 indicate that supervised fine-tuning leads to more accurate internal beliefs.
@@ B.7 Supervised Fine-tuning Results
Prec. Rec. Prec. Rec.
@@
Figure 8: Layer-wise AUROC of belief scores in Qwen2.5-7B (Base) and its fine-tuned variants (SFT) on Wikidata.
Figure 9: Layer-wise AUROC of belief scores in Olmo2-7B (Base) and its fine-tuned variants (SFT) on Wikidata.
B.8 Generalization to Math Reasoning
Although our analysis primarily focuses on factoid QA, we further investigate whether belief steering extends to math reasoning, using the GSM8k dataset (Cobbe et al., 2021).
We first collect Llama3.1-8B’s trajectories that produce incorrect final answer via greedy decoding, and use GPT-4.1 to annotate the first incorrect token. This results in a total of 188 examples (where $\text{accuracy}=0$ with no steering by construction). We then apply negative belief steering, without altering any other generation setting, and evaluate accuracy on this subset. To moderately amplify the effect, we steer layers 0-18 and intervene on the first incorrect token plus its preceding nine tokens (a randomly selected hyperparameter). Since math reasoning depends on multi-step computation, modifying early hidden states can propagate and influence later inference. We also report positive belief steering to reveal unintended effects. Results are presented in Table 21, and a case study is illustrated below.
Accuracy
Belief- Steering 37.77%
Belief+ Steering 17.55%
Table 21: Steering results on GSM8k.
Recall that the belief vector is derived from a quite different dataset UTQA. Negative belief steering still activates retraction behavior in math reasoning and improve the final accuracy by around 20%, demonstrating the robustness and generalization of our interpretability findings.
Case Study [QUESTION]
Carla is downloading a 200 GB file. Normally she can download 2 GB/minute, but 40% of the way through the download, Windows forces a restart to install updates, which takes 20 minutes. Then Carla has to restart the download from the beginning. How long does it take to download the file?
[RESPONSE] – No Steering ✗
… 3. After the restart, Carla has to download the remaining 60% of the file (100% - 40% = 60%)…
[RESPONSE] – Belief- Steering ✓
… 3. After the restart, Carla has to download the remaining 60% [^7] Steering is applied at this token and the preceding nine tokens. of the file, but Carla has to restart the download from the beginning. So Carla has to download the entire 200 GB again…
[RESPONSE] – Belief+ Steering ✗
… 3. After the restart, Carla has to download the remaining 60% .
  60% of 200 GB is 0.6 * 200 = 120 GB…